# An Objective Improvement Approach

### v6

@@ 1. Introduction
Solving graph games is the central and most expensive step in many model checking [Koz83, EJS93, Wil01, dAHM01, AHK02, SF06a], satisfiability checking [Koz83, Var98, Wil01], and synthesis [Pit06, SF06b] algorithms. Progress in algorithms for solving graph games will, therefore, allow for the development of more efficient model checkers and contribute to bringing synthesis techniques to practice.
There is a hierarchy among the graph games mentioned earlier, with simple and well-known reductions from parity games to mean payoff games, from mean-payoff games to discounted payoff games, and from discounted payoff games to simple stochastic games such as the ones from [ZP96], while no reductions are known in the other direction. Therefore, one can solve instances of all these games by using an algorithm for stochastic games. All of these games are in UP and co-UP [Jur98], while no tractable algorithm is known.
Most research has focused on parity games: as the most special class of games, algorithms have the option to use the special structure of their problems, and they are most directly linked to the synthesis and verification problems mentioned earlier. Parity games have thus enjoyed a special status among graph games and the quest for efficient algorithms [EL86, EJ91, McN93, ZP96, BCJ+97, Zie98, Obd03, BDM16, BDM18] to solve them has been an active field of research during the last decades, which has received further boost with the arrival of quasi-polynomial techniques [JL17, FJdK+19, LB20, LPSW20, DS22, CJK+22, BDM+24].
Interestingly, the only class of efficient techniques for solving parity games that does not (yet) have a quasi-polynomial approach is strategy improvement algorithms [Lud95, Pur95, VJ00, BV07, Sch08, Fea10, STV15], a class of algorithms closely related to the Simplex for linear programming, known to perform well in practice. Most of these algorithms reduce to mean [BV07, Sch08, STV15, BDM20] or discounted [Lud95, Pur95, FGO20, Koz21] payoff games.
With the exception of the case in which the fixed-point of discounted payoff games is explicitly computed [ZP96], all these algorithms share a disappointing feature: they are inherently non-symmetric approaches for solving an inherently symmetric problem. However, some of these approaches have a degree of symmetry. Recursive approaches treat even and odd colours symmetrically, one at a time, but they treat the two players very differently for a given colour. Symmetric strategy improvement [STV15] runs a strategy improvement algorithms for both players in parallel, using the intermediate results of each of them to inform the updates of the other, but at heart, these are still two intertwined strategy improvement algorithms that, individually, are not symmetric. This is due to the fact that applying strategy improvement itself symmetrically can lead to cycles [Con93].
@@ 2. Preliminaries
A pair of strategies $\sigma_{\min}$ and $\sigma_{\max}$, one for each player, defines a unique play $\rho(v,\sigma_{\min},\sigma_{\max})$ from each vertex $v\in V$. Discounted payoff games are positionally determined [ZP96]:
@@ 3. Outline and Motivation Example
The optimal decision of the maximiser is, therefore, to move on the first time, which yields the maximal reward of $\frac{\lambda}{1-\lambda}$. Every vertex $v$ has some outgoing edge(s) $e=(v,v^{\prime})$ where $\mathsf{val}_{\mathcal{G}}(v)=w_{e}+\lambda_{e}\mathsf{val}_{\mathcal{G}}(v^{\prime})$ holds [ZP96]; these edges correspond to the optimal decisions for the respective player.
@@ 5.3. Games with Efficient Objective Improvement
As our definition of sharp games simply says that every basic solution is nondegenerate, every step of a simplex algorithm (outside of optimal solutions) will provide a strict improvement [BT97, Theorem 3.3], so that $\mathcal{G}$ is $\sigma$-improving for every joined strategy $\sigma$ and all solutions $\mathsf{val}$ defined by bases that are not optimal w.r.t. $f_{\sigma}$.
@@ 5.5. Mixing Pivoting on the Simplex and of the Objective
The other two updates, however, can be freely mixed, as they both bring down the value of $f_{\sigma}(\mathsf{val})$ by applying local changes. That the improvements from (1) are given preference in the algorithm is a choice made to keep the implementation of the algorithm for using linear programs open, allowing, for example, to use ellipsoid methods [Kha79] or inner point methods [Kar84] to keep this step tractable.
@@ 7. Discussion
There is widespread belief that mean payoff and discounted payoff games have two types of algorithmic solutions: value iteration [FGO20, Koz21] and strategy improvement [Lud95, Pur95, BV07, Sch08, STV15]. We have added a third method, which is structurally different and opens a new class of algorithms to attack these games. Moreover, our new symmetric approach has the same proximity to linear programming as strategy improvement algorithms, which is an indicator of efficiency.
@@
A second question is whether this method as whole can be turned into an inner point method [Kar84]. If so, this could be a first step towards showing tractability of discounted payoff games – which would immediately extend to mean-payoff and parity games.

### v7

@@ 1. Introduction
Solving graph games is the central and most expensive step in many model checking [26, 15, 40, 12, 1, 34], satisfiability checking [26, 38, 40], and synthesis [32, 35] algorithms. Progress in algorithms for solving graph games will, therefore, allow for the development of more efficient model checkers and contribute to bringing synthesis techniques to practice.
There is a hierarchy among the graph games mentioned earlier, with simple and well-known reductions from parity games to mean payoff games, from mean-payoff games to discounted payoff games, and from discounted payoff games to simple stochastic games such as the ones from [42], while no reductions are known in the other direction. Therefore, one can solve instances of all these games by using an algorithm for stochastic games. All of these games are in UP and co-UP [22], while no tractable algorithm is known.
Most research has focused on parity games: as the most special class of games, algorithms have the option to use the special structure of their problems, and they are most directly linked to the synthesis and verification problems mentioned earlier. Parity games have thus enjoyed a special status among graph games and the quest for efficient algorithms [17, 16, 30, 42, 9, 41, 31, 3, 5, 4] to solve them has been an active field of research during the last decades, which has received further boost with the arrival of quasi-polynomial techniques [21, 18, 27, 28, 14, 10, 2].
Interestingly, the only class of efficient techniques for solving parity games that does not (yet) have a quasi-polynomial approach is strategy improvement algorithms [29, 33, 39, 8, 37, 19, 36], a class of algorithms closely related to the Simplex for linear programming, known to perform well in practice. Most of these algorithms reduce to mean [8, 37, 36, 6] or discounted [29, 33, 20, 25] payoff games.
With the exception of the case in which the fixed-point of discounted payoff games is explicitly computed [42], all these algorithms share a disappointing feature: they are inherently non-symmetric approaches for solving an inherently symmetric problem. However, some of these approaches have a degree of symmetry. Recursive approaches treat even and odd colours symmetrically, one at a time, but they treat the two players very differently for a given colour. Symmetric strategy improvement [36] runs a strategy improvement algorithms for both players in parallel, using the intermediate results of each of them to inform the updates of the other, but at heart, these are still two intertwined strategy improvement algorithms that, individually, are not symmetric. This is due to the fact that applying strategy improvement itself symmetrically can lead to cycles [11].
@@ 2. Preliminaries
A pair of strategies $\sigma_{\min}$ and $\sigma_{\max}$, one for each player Min and Max, defines a unique play $\rho(v,\sigma_{\min},\sigma_{\max})$ from each vertex $v\in V$. Discounted payoff games are positionally determined [42]:
@@ 3. Outline and Motivation Example
The optimal decision of the maximiser is, therefore, to move on the first time, which yields the maximal reward of $\frac{\lambda}{1-\lambda}$. Every vertex $v$ has some outgoing edge(s) $e=(v,v^{\prime})$ where $\mathsf{val}_{\mathcal{G}}(v)=w_{e}+\lambda_{e}\mathsf{val}_{\mathcal{G}}(v^{\prime})$ holds [42]; these edges correspond to the optimal decisions for the respective player.
@@ 5.3. Games with Efficient Objective Improvement
As our definition of sharp games simply says that every basic solution is nondegenerate, every step of a simplex algorithm (outside of optimal solutions) will provide a strict improvement [7, Theorem 3.3], so that $\mathcal{G}$ is $\sigma$-improving for every joined strategy $\sigma$ and all solutions $\mathsf{val}$ defined by bases that are not optimal w.r.t. $f_{\sigma}$.
@@ 5.5. Mixing Pivoting on the Simplex and of the Objective
The other two updates, however, can be freely mixed, as they both bring down the value of $f_{\sigma}(\mathsf{val})$ by applying local changes. That the improvements from (1) are given preference in the algorithm is a choice made to keep the implementation of the algorithm for using linear programs open, allowing, for example, to use ellipsoid methods [24] or inner point methods [23] to keep this step tractable.
@@ 7. Discussion
There is widespread belief that mean payoff and discounted payoff games have two types of algorithmic solutions: value iteration [20, 25] and strategy improvement [29, 33, 8, 37, 36]. We have added a third method, which is structurally different and opens a new class of algorithms to attack these games. Moreover, our new symmetric approach has the same proximity to linear programming as strategy improvement algorithms, which is an indicator of efficiency.
@@
A second question is whether this method as whole can be turned into an inner point method [23]. If so, this could be a first step towards showing tractability of discounted payoff games – which would immediately extend to mean-payoff and parity games.